The initial price for a stadium is $800,000,000. There will be a 5% adjustment to the price, a 3% discount to the price, and $85,000,000 of revenue from the sale of previous equipment and land. If the projected future cash flow is $675,000,000. What is the Net Present Value of this project.?
Answer #1

The formula for calculating Net Present Value is,

NPV = Present value of cash inflows - initial investment

Calculation of initial investment

Initial price = 800,000,000

It is mentioned that there will be 5% adjustment to the price, so price after adjusting 5%

800,000,000 - (800,000,000 * 5%) = 760,000,000

There is also 3% discount to the price. So price after providing discount is,

760,000,000 - (800,000,000 * 3%) = 736,000,000

Revenue from the sale of previous equipment and land = 85,000,000

It should be deducted from the initial price for calculating Initail cash outflow,

736,000,000 - 85,000,000 = 651,000,000

Therefore initial investment = 651,000,000

Calculation of NPV,

Projected future cash flow = 675,000,000

Initial investment = 651,000,000

NPV = 675,000,000 - 651,000,000 = 24,000,000

The decision criteria for NPV is, If the NPV is positive (greater than 0) the project proposal can be accepted and if NPV is negative the project should be rejected. Here the project has NPV greater than 0 and thus it is acceptable.

Note: As per the rule, Present value of cash flows are taken in NPV calculation ie, discounted cash flows. In this question discount rate is not mentioned and projected cash flow from the project is given as 675,000,000. Therefore assume 675,000,000 is the present value of cash inflows of the project.